OpenStudy (ttp):

Evaluate the variable a raised to the second power times the variable a raised to the negative second power for a equals 8 for a = 8. 1 negative one sixty-fourth -64 64

OpenStudy (anonymous):

\[8^2 \times 8^{-2} = 64 \times \frac{1}{64}\] Remember: 8^-2 = 1/ 8^2

OpenStudy (ttp):

So B?

OpenStudy (anonymous):

multiply 1/64 with 64 and get

OpenStudy (anonymous):